the perimeter of a rectangle is 84 inches, the lenght exceeds the width by 24 inches finf the lenght and width
Answer by mananth(16946)

Perimeter = 84
Length (x+24)
Width =x ft

Perimeter = 2(x+24 +x)
Perimeter = 2(2x+24)

84 = 4 x + 48
84 - 48 = 4 x
84 -48 = 4 x
36 = 4 x
x= 9 ft Width
Length = 33 ft